I need to find the equation of this line. The answer should be written in slope intercept form. Please show all steps.
The line is perpendicular to x -5y = 4 and passes through
the point (-1, 1)

Found 2 solutions by drk, dabanfield:
Answer by drk(1908)

Step 1: write the equation in y = mx + b.
Y = (1/5)X -4/5.
step 2: find the slope of the line from step 1.
the slope = 1/5
step 3: perpendicular slope means take the negative flip of the slope from step 2. So 1/5 become -5. This is the slope we need.
step 4: put new slope from step 3 and a point into y = mx + b to find b.
1 = (-1)(-5) + b
b = -4.
step 5: write the answer.
Y = -5x -4.

Answer by dabanfield(803)

The line is perpendicular to x -5y = 4 and passes through
the point (-1, 1)
The slope-intercept form of the required line is:
y = m*x +b
The slope-intercept form of the given perpendicular line is:
y = (1/5)*x - 4/5
The slope of this line is 1/5. The line perpenciculor to this would have a slope that was the negative reciprocal of 1/5 or -5
So the slope of the required line is -5 and we have:
y = -5*x + b
Since this line passes through the point (-1,1) we know that:
1 = -5*-1 + b
b = -4
The equation of the second line is then:
y = -5*x -4